OpenStudy (anonymous):

What is the solution of the system? y = x - 14 y = -4x + 11 Explain, please. Medal for best answer.

OpenStudy (anonymous):

What are you solving for? and how?

OpenStudy (anonymous):

Try elimination. Multiply the entire top equation by 4 and add to the 2nd equation. 4y ... 4x - 56 + y ... -4x + 11 The x's cancel, solve for y, then solve for x.

OpenStudy (anonymous):

5y = (-56+11)

OpenStudy (anonymous):

So, 5y = -45?

OpenStudy (anonymous):

Divide both sides by 5 to find y. Then plug in the y-value into y = x - 14, to solve for x

OpenStudy (anonymous):

y = -9 Then, -9 = x -14 is the equation I work with?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you can also wt=rite is as x - 14 = -9

OpenStudy (anonymous):

Would I subtract both sides by -9? I'm a bit confused.

OpenStudy (anonymous):

Or, add 9 to both sides?

OpenStudy (anonymous):

x - 14 = -9 solve for x +14 +14 x -14 + 14 = - 9 +14 x = -9 +14 x = ?

OpenStudy (anonymous):

Oh, I was trying to use the wrong side it seems. x = 5?

OpenStudy (anonymous):

The solution set is (5,-9)

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Great! Thanks for all the help, hopefully in the future I'll be able to figure it all out easier.

OpenStudy (anonymous):

Now, to close the question... thanks again!